Suppose the production function a firm producing Q requires the use of only labor and


capital. The production function for the product is given by: Q-1/4K3/4. The amount of capital used is fixed at 5 units. Then,


A. Determine MPL and APL function


B. Determine MP and AP function

C. Compute MRTSKL & MRTS LK


1
Expert's answer
2022-01-10T09:59:11-0500

Solution:

A.). Q = L1/4K3/4

MPL = "\frac{\partial Q} {\partial L}" = 0.25L-0.25K0.75


APL = "\frac{ Q} {L}" = 0.25K0.75 "\div" L         

 

B.). MPK = "\frac{\partial Q} {\partial K}" = 0.75L0.25K-0.75


APK = "\frac{ Q} {K}" = 0.25L0.25 "\div" K        

 

C.). MRTSKL =  MPL "\div" MPK = 0.25L-0.25K0.75 "\div" 0.75L0.25K-0.75 = 0.3K1.5 "\div" L0.5


MRTSLK =  MPK "\div" MPL = 0.75L0.25K-0.75 "\div" 0.25L-0.25K0.75 = 3L0.5 "\div" K1.5